OpenStudy (anonymous):

If the length of each side of a cube is halved, is the surface area halved? Use an example to support your answer.

OpenStudy (anonymous):

Please help ;)

OpenStudy (anonymous):

The answer is no you can use x=4 and y=6 for a nice example

OpenStudy (anonymous):

@witcon Thank you!

OpenStudy (anonymous):

since x*6=24, what is 2*3 etc.

OpenStudy (anonymous):

I dont get it. sorry i thought i did.

OpenStudy (anonymous):

hah I misread it as square not cube, dumb me, anyway we can use the same technique to show it

OpenStudy (anonymous):

@witcon lol.

OpenStudy (anonymous):

if we set the sides of a cube as 4 units, then the surface area is 4*4*6=96

OpenStudy (anonymous):

units squared

OpenStudy (anonymous):

if we halve the side lengths then its 2*2*6=24=/=96/2

OpenStudy (anonymous):

I need this answered in like 1 minute. so do you think you can hurry please?

OpenStudy (anonymous):

6x^2, now put x as x/2, see what happens.

OpenStudy (anonymous):